# PainDiffusion: Learning to Express Pain

### v2

@@ I INTRODUCTION
In this work, we focus on synthesizing facial reactions based on embodiment signals. We exclusively use the BioVid HeatPain Database [11], which is the only dataset that includes direct recordings of stimuli, physiological signals, and naturalistic facial reactions. This restricts our ability to generalize across diverse patient populations and cultural backgrounds. Our contributions are threefold: (1) We introduce PainDiffusion, a model designed to generate pain-related facial expressions with arbitrary-length predictions, making it suitable for robotic applications. It incorporates intrinsic characteristics such as expressiveness and emotion, allowing for more controllable and personalized generation. (2) We propose a new set of baselines and metrics to effectively evaluate the quality and accuracy of pain expressions generated by our model. (3) Finally, we integrate PainDiffusion with a robotic elbow and have rehabilitation clinicians assess the pain reactions to their actions.
@@ II RELATED WORKS
In training robots, earlier work in pain synthesizing primarily focused on recognizing pain situations and selecting expressions from a predefined set [10, 11]. Among pain-related research, most efforts have focused on recognition and classification using either traditional machine learning or deep learning techniques [12, 9, 6]. However, those methods often results in expressions that are unnatural and not automatic. More recently, research has shifted towards using generative models for facial expressions [13]. Meanwhile, diffusion models and score-based models have emerged as powerful tools for generating images and videos. These models have now achieved state-of-the-art performance, particularly in generating realistic images and videos [14, 15, 16]. In human action generation, recent work is increasingly focusing on diffusion models to achieve state-of-the-art results [17, 18, 19].
@@ III-E Diffusion Forcing
As diffusion forcing requires, we train our model with random noise levels for each temporal frame. During sampling, we apply a scheduling matrix that denoises a window of frames $w$ and then shifts the window by a horizon step $h$, ensuring there are some overlapping context frames $w-h>0$. This overlap determines how quickly the model can respond to oncoming stimuli signals and how many past frames affect to the current generation. For example, with a small forward time, a sampling rate of 32Hz, and $h=16$, the model would react with a delay of 0.5 seconds and it consider 16 frames as context frames.
@@ IV-B Experiments setup
Baselines. We establish three baselines to validate the model’s effectiveness. Because generating :
@@
The experiment is implemented using the jsPsych framework222jsPsych is a JavaScript framework for creating behavioral experiments that run in a web browser.. From a dataset of 50 random validation video samples, each question is randomly selected from the dataset. To assess diversity, we generate four variations of the same stimuli. For both evaluating temporal consistency and diversity, participants rate the strength of the reaction on a 5-point scale. Each question is separated by a fixed screen in one second to inform the user of the boundary between the questions. We learn that cropping mouth and eye regions helps users better assess diversity as it limits the cognitive load of the comparison. We have 2/3 of diversity questions in the format of cropping eye and mouth.
@@ IV-C Qualitative results
We employed the Farneback method [39] to detect motion, evaluating the magnitude and area of facial movements in the generated sequences. To assess controllability, we varied the stimuli intensity, emotional configurations, and expressiveness settings. The mean movements are presented in Fig. 8. Our findings indicate that stronger stimuli levels elicit more pronounced facial motions. Neutral emotional states correspond to minimal facial movement, whereas sadness and happiness lead to more extensive movements in response to pain stimuli. Interestingly, higher expressiveness settings result in less intense movements.
@@ IV-D Quantitative results
In terms of expression diversity, PainDiffusion generates facial expressions that are not only closer to the ground truth but also exhibit greater variation across multiple predictions (lower PainDist, higher PainDivrs). Additionally, during a prediction sequence, our model achieves the highest expression variance, as indicated by higher PainVar. Despite this higher variance, analysis using the Farneback method [39] shows that PainDiffusion’s facial movements are more concentrated in key areas such as the mouth, eyebrows, chin, and nose (Fig. 9), which closely mirrors the ground truth.
@@ V CONCLUSION
In this work, we presented PainDiffusion, a model designed to generate appropriate facial expressions in response to pain stimuli, with the ability to control pain expressiveness characteristics. PainDiffusion leverages diffusion forcing within a latent diffusion model that captures temporal information, enabling it to generate long-term predictions efficiently, making it suitable for robotic applications. The model generates more diverse and concentrated expressions compared to the autoregressive baseline, approaching the random baseline in terms of diversity while outperforming all baselines in terms of pain PSPI similarity and correlation.
The current approach focuses exclusively on non-verbal expressions, with controllability limited to emotional state and pain expressiveness. Future work could extend this framework by incorporating additional factors, such as material wear characteristics or real-time physical damage, to enhance the accuracy of pain expression synthesis. Integrating auditory cues, such as vocal reactions and breathing patterns, could further improve realism by enabling a more natural and multimodal representation of pain. Additionally, the availability of a more diverse and naturalistic facial pain dataset would be beneficial in improving the generalization of automatic synthesis methods across varied demographics and clinical contexts.

### v3

@@ I INTRODUCTION
In this work, we focus on synthesizing facial reactions based on embodiment signals. We exclusively use the BioVid HeatPain Database [11], which is the only dataset that includes direct recordings of stimuli, physiological signals, and naturalistic facial reactions. This restricts our ability to generalize across diverse patient populations and cultural backgrounds. Due to limited resources, we can currently only deploy the model on virtual avatars, reserving real-world deployment for future work. Our contributions are threefold: (1) We introduce PainDiffusion, a model designed to generate pain-related facial expressions with arbitrary-length predictions, making it suitable for robotic applications. It incorporates intrinsic characteristics such as expressiveness and emotion, allowing for more controllable and personalized generation. (2) We propose a new set of baselines and metrics to effectively evaluate the quality and accuracy of pain expressions generated by our model. (3) Finally, we integrate PainDiffusion with a robotic elbow and have rehabilitation clinicians assess the pain reactions to their actions.
@@ II RELATED WORKS
In training robots, earlier work in pain synthesizing primarily focused on recognizing pain situations and selecting expressions from a predefined set [10, 12]. Among pain-related research, most efforts have focused on recognition and classification using either traditional machine learning or deep learning techniques [13, 9, 6]. However, those methods often result in expressions that are unnatural and not automatic. More recently, research has shifted towards using VQ-VAE and autoregressive models for facial expressions [14]; however, their decoders can produce jittery motion at the boundaries of the motion sequence. In human action generation, recent work is increasingly focusing on diffusion models to achieve state-of-the-art results [15, 16, 17], which generate motions in a seamless way. However, these models typically assume fixed-length window generation and require many forward passes, making them unsuitable for robotic applications. Therefore, to address the lack of naturalness in synthesized pain expressions, this paper proposes a novel approach using diffusion models to generate realistic, seamless, and automatic pain behaviors for robots.
@@ III-E Diffusion Forcing
As diffusion forcing requires, we train our model with random noise levels for each temporal frame. During sampling, we apply a scheduling matrix that denoises a window of frames $w$ and then shifts the window by a horizon step $h$, ensuring there are some overlapping context frames wh>0w-h>0. This overlap determines how quickly the model can respond to oncoming stimuli signals and how many past frames affect to the current generation. For example, with a small forward time, a sampling rate of 32 Hz, and $h=16$, the model would react with a delay of 0.5 seconds and it consider 16 frames as context frames.
@@ IV-B Experiments setup
Baselines. We establish three baselines to validate the model’s effectiveness:
@@
The experiment is implemented using the jsPsych framework222jsPsych is a behavioral experiments framework that run in a web browser.. From a dataset of 50 random validation video samples, each question is randomly selected from the dataset. To assess diversity, we generate four variations of the same stimuli. For both evaluating temporal consistency and diversity, participants rate the strength of the reaction on a 5-point scale. Each question is separated by a fixed screen in one second to inform the user of the boundary between the questions. We learn that cropping mouth and eye regions helps users better assess diversity as it limits the cognitive load of the comparison. We have 2/3 of diversity questions in the format of cropping eye and mouth.
@@ IV-C Qualitative results
We employed the Farneback method [39] to analyze facial motion patterns, quantifying both the magnitude and distribution of movements in the generated sequences. To evaluate controllability, we systematically varied three parameters: stimuli intensity, emotional state, and expressiveness settings. The resulting mean movements are visualized in Fig. 8. As expected, increasing stimuli levels consistently produced more pronounced facial movements. Emotional states significantly modulated these responses - neutral states showed minimal movement, while sadness and happiness amplified pain-related facial motions. Notably, higher expressiveness settings (measured by average PSPI [34] scores) correlated with reduced movement intensity. This counterintuitive finding suggests that individuals who maintain higher baseline pain expressions in the dataset may actually exhibit more subtle changes in facial movement, possibly indicating a different strategy for communicating pain intensity.
@@ IV-D Quantitative results
For expression quality, PainDiffusion generates facial expressions that maintain better fidelity to ground truth while exhibiting greater diversity across multiple predictions, as indicated by lower Dist and higher Divrs metrics. Within individual sequences, our model achieves the highest expression variance (Var). Analysis using the Farneback method [39] reveals that PainDiffusion concentrates facial movements in key areas - mouth, eyebrows, chin, and nose - closely matching ground truth patterns (Fig. 9).
In terms of real-time robotic applications, diffusion forcing proves particularly effective. While it shows slightly lower similarity scores and reduced diversity metrics (Divrs, Var) compared to full sequence diffusion, it achieves significantly better correlation and distance metrics. This makes diffusion forcing the more suitable choice for real-time robotic implementations where stable long-sequence generation is critical.
@@ V CONCLUSION
In this work, we introduced PainDiffusion, a novel model for generating appropriate and controllable facial expressions in response to pain stimuli. By leveraging diffusion forcing within a latent diffusion model, PainDiffusion effectively captures temporal dynamics to produce efficient, long-term predictions suitable for robotic applications. Our evaluation demonstrates that the model generates more diverse and specific expressions than baselines, achieving superior performance in pain-specific metrics like PSPI similarity and correlation.
Despite promising results, several areas for future work are crucial. The current model is limited to non-verbal expressions, with control restricted to emotional state and expressiveness. Furthermore, testing was only conducted on a virtual avatar, emphasizing the need for evaluation on a physical robot interface. Crucially, qualitative insights from clinicians highlight the necessity of culturally-aware, personalized medical technology. This finding points to a significant limitation of current datasets and the need for more diverse and naturalistic facial pain data.